# Why are (almost) all the protostellar outflows aligned in Serpens Main?

### v2

@@ 3.2 Identifying outflows
1 | 277.45017 | 1.27892 | 0.39 | 1.12 | A | SMM 9 (SH 2-68N) | ||
2 | 277.45025 | 1.26917 | 0.59 | 1.11 | B | SMM 1 (S7) | ||
3 | 277.45296 | 1.28233 | 0.91 | 1.23 | A | S68Ncc | ||
4 | 277.45471 | 1.27225 | 0.9-2.8 | 1.18 | A | OO Ser | ||
5 | 277.45521 | 1.275431 | 1.4 | 1.06 | A | EC37 (V370 Ser) | ||
6 | 277.45663 | 1.28506 | 1.3 | – | A | S68Nb2 | ||
7 | 277.45704 | 1.24914 | 0.51-1.7 | – | A | SMM 1b | ||
8 | 277.45742 | 1.25581 | 1.5-5.1 | 1.03 | A | SMM 1a | ||
9 | 277.45946 | 1.23919 | 0.89-1.4 | 1.04 | A | SSTc2d J182950.5+01141 | ||
10 | 277.46321 | 1.27800 | 1.7 | – | A | EC 53 | ||
11 | 277.46742 | 1.26347 | 1.3 | 1.29 | A | Serp 20 | ||
12 | 277.46833 | 1.25169 | 1.2 | 1.02 | A | No identification | ||
13 | 277.47400 | 1.22158 | 0.59 | 5.56 | B | No identification | ||
14 | 277.47996 | 1.22283 | 0.63 | – | A | Serpens 56 | ||
15 | 277.49504 | 1.24622 | 0.33 | – | B | No identification | ||
16 | 277.49642 | 1.23522 | 1.4 | 1.32 | A | SMM 3 | ||
17 | 277.49646 | 1.21064 | 1.2 | – | A | Serpens 9 | ||
18 | 277.50167 | 1.19583 | 1.1 | 1.02 | B | SMM 11 | ||
19 | 277.50296 | 1.21603 | 1.7 | – | C | Ser-emb 4E | ||
20 | 277.51067 | 1.24542 | 1.7 | – | A | 2MASS J18300491+0114393 | ||
21 | 277.48688 | 1.24633 | – | – | A | [EC92] 82 | ||
22 | 277.50621 | 1.25431 | – | – | A | Shd 2 |

### v3

@@ 3.2 Identifying outflows
| 1 | 277.45017 | 1.27892 | 0.39 | 1.12 | A | SMM 9 (SH 2-68N) | ||
| 2 | 277.45025 | 1.26917 | 0.59 | 1.11 | B | SMM 1 (S7) | ||
| 3 | 277.45296 | 1.28233 | 0.91 | 1.23 | A | S68Ncc | ||
| 4 | 277.45471 | 1.27225 | 0.9-2.8 | 1.18 | A | OO Ser | ||
| 5 | 277.45521 | 1.275431 | 1.4 | 1.06 | A | EC37 (V370 Ser) | ||
| 6 | 277.45663 | 1.28506 | 1.3 | – | A | S68Nb2 | ||
| 7 | 277.45704 | 1.24914 | 0.51-1.7 | – | A | SMM 1b | ||
| 8 | 277.45742 | 1.25581 | 1.5-5.1 | 1.03 | A | SMM 1a | ||
| 9 | 277.45946 | 1.23919 | 0.89-1.4 | 1.04 | A | SSTc2d J182950.5+01141 | ||
| 10 | 277.46321 | 1.27800 | 1.7 | – | A | EC 53 | ||
| 11 | 277.46742 | 1.26347 | 1.3 | 1.29 | A | Serp 20 | ||
| 12 | 277.46833 | 1.25169 | 1.2 | 1.02 | A | No identification | ||
| 13 | 277.47400 | 1.22158 | 0.59 | 5.56 | B | No identification | ||
| 14 | 277.47996 | 1.22283 | 0.63 | – | A | Serpens 56 | ||
| 15 | 277.49504 | 1.24622 | 0.33 | – | B | No identification | ||
| 16 | 277.49642 | 1.23522 | 1.4 | 1.32 | A | SMM 3 | ||
| 17 | 277.49646 | 1.21064 | 1.2 | – | A | Serpens 9 | ||
| 18 | 277.50167 | 1.19583 | 1.1 | 1.02 | B | SMM 11 | ||
| 19 | 277.50296 | 1.21603 | 1.7 | – | C | Ser-emb 4E | ||
| 20 | 277.51067 | 1.24542 | 1.7 | – | A | 2MASS J18300491+0114393 | ||
| 21 | 277.48688 | 1.24633 | – | – | A | [EC92] 82 | ||
| 22 | 277.50621 | 1.25431 | – | – | A | Shd 2 |